# Diff-Reg v1: Diffusion Matching Model for Registration Problem

### v3

@@ IV Experiments
We provide a visualization to demonstrate our method’s effectiveness in Fig.4. For a fair comparison, we exploit the “metric index” (i.e., the test point set in the 4DMatch/4DLoMatch dataset, in blue color) of the source point cloud for all methods. Taking the predicted correspondences from RoITr [2], GeoTr [2], and “Our (Diffusion/20)” as anchor correspondences, we calculate the deformation flow for the source test points by applying neighborhood k-nearest neighbors (knn) interpolation based on the anchors. The deformable registration of the bear’s two front paws in the first row and fourth column reveals that dealing with ambiguous matching patches of asymmetric objects can be highly challenging. However, our denoising process is capable of handling this scenario perfectly. The deformable registration results from the top three rows (from the 4DMatch benchmark) indicate that the baseline methods struggle to compute reliable and consistent correspondences between scans with large deformations. The bottom two rows (from the 4DLoMatch benchmark) also demonstrate that low overlapping combined with deformation results in a disaster. These visualizations demonstrate that our denoising module in the matching matrix space may provide a more effective approach for tackling these challenges. We provide more additional visualizations about tough cases in Fig.7. The top three lines (in Fig.7) are from 4DMatch, while the bottom three lines are from the 4DLoMatch dataset. The results in the bottom three lines demonstrate that our denoising module can effectively handle scenes with ”large deformations + low overlap,” whereas the top three lines show that our method excels in registering asymmetric objects.
IV-A3 Rigid 3DMatch/3DLoMatch benchmark.
Datasets.
3DMatch [31] is an indoor benchmark for 3D matching and registration. Following [11, 6, 8], we split it to 46/8/8 scenes for training/validation/testing. The overlap ratio between scan pairs in 3DMatch/3DLoMatch is about $>30\%$/$10\%-30\%$.
Metrics.
Following [11, 6, 66], we utilize three evaluation metrics to evaluate our method and other baselines: (1) Inlier Ratio (IR): The proportion of accurate correspondences in which the distance falls below a threshold (i.e., $0.1m$) based on the ground truth transformation. (2) Feature Matching Recall (FMR): The percentage of matched pairs that have an inlier ratio exceeding a specified threshold (i.e., $5\%$). (3) Registration Recall (RR): The fraction of successfully registered point cloud pairs with a predicted transformation error below a certain threshold (e.g., RMSE $<$ 0.2).
Results.
We compared our method with some state-of-the-art feature matching based methods: FCGF [65], D3Feat [3], Predator [11], Lepard [8], GeoTr [6], and RoITr [2]. Diff-Reg(diffusion/1) and Diff-Reg(diffusion/20) represent our denoising module with one or twenty reverse sampling steps, respectively. As demonstrated in Table III, our method Diff-Reg(diffusion/20) achieves the highest registration recall on the 3DMatch benchmark.
@@ Results.
TABLE IV: Evaluation results on RGB-D Scenes V2 [4]. The best results are highlighted in bold, and the second-best results are underlined.
Model Scene-11 Scene-12 Scene-13 Scene-14 Mean
@@
[…]
FCGF-2D3D [65] 6.8 8.5 11.8 5.4 8.1
P2-Net [65] 9.7 12.8 17.0 9.3 12.2
Predator-2D3D [11] 17.7 19.4 17.2 8.4 15.7
2D3D-MATR [4] 32.8 34.4 39.2 23.3 32.4
FreeReg [38] 36.6 34.5 34.2 18.2 30.9
Diff-Reg(dino) 38.6 37.4 45.4 31.6 38.3
Diff-Reg(dino/backbone) 44.9 49.5 38.3 33.1 41.4
Diff-Reg(dino/diffusion/1) 47.5 48.9 32.8 22.4 37.9
Diff-Reg(dino/diffusion/10) 47.2 48.7 32.9 22.4 37.8
[…]
FCGF-2D3D [65] 11.10 30.40 51.50 15.50 27.10
P2-Net [65] 48.60 65.70 82.50 41.6 59.60
Predator-2D3D [11] 86.10 89.20 63.90 24.30 65.90
2D3D-MATR [4] 98.60 98.00 88.70 77.90 90.80
FreeReg [38] 91.90 93.40 93.10 49.60 82.00
Diff-Reg(dino) 100. 100. 89.70 81.9 92.9
Diff-Reg(dino/backbone) 100. 100. 92.8 91.2 96.0
Diff-Reg(dino/diffusion/1) 100. 100. 88.7 76.5 91.3
Diff-Reg(dino/diffusion/10) 100. 100. 88.7 77.0 91.4
[…]
@@
Diff-Reg(dino/diffusion/1) 98.6 100. 87.6 66.8 88.3
Diff-Reg(dino/diffusion/10) 98.6 96.1 83.5 63.7 85.5
@@
IV-B 2D3D Registration Task: RGB-D SCENES V2
2D3D registration task is non-trivial because the 2D image data is a perspective projection of the 3D scene, which creates the scale ambiguity problem[4]. Since there is no good robust differentiable PnP[67] solver that can be integrated into our denoising module, we deploy the SOTA depth estimation model Depth anything[52] to generate “affine-invariant depth” for the image. Subsequently, we utilize the camera’s intrinsic matrix to project a point cloud onto a new non-metric point cloud, which is paired with the corresponding real point cloud in the dataset, creating a challenge scale ambiguous registration problem. To alleviate the scale ambiguity in the image data, we utilize the SOTA self-supervised foundation model DINOv2 [39] to enhance the image features.
TABLE V: Evaluation results on RGB-D Scenes V2[4]. The best results are highlighted in bold, and the second-best results are underlined.
Model Scene-11 Scene-12 Scene-13 Scene-14 Mean
Mean depth (m) 1.74 1.66 1.18 1.39 1.49
[…]
FCGF-2D3D[65] 26.4 41.2 37.1 16.8 30.4
P2-Net[37] 40.3 40.2 41.2 31.9 38.4
Predator-2D3D[11] 44.4 41.2 21.6 13.7 30.2
2D3D-MATR[4] 63.9 53.9 58.8 49.1 56.4
FreeReg+Kabsch[38] 38.7 51.6 30.7 15.5 34.1
FreeReg+PnP[38] 74.2 72.5 54.5 27.9 57.3
Diff-Reg(dino) 87.5 86.3 63.9 60.6 74.6
Diff-Reg(dino/backbone) 79.2 86.3 75.3 71.2 78.0
Diff-Reg(dino/diffusion/1) 94.4 98.0 85.6 63.7 85.4
Diff-Reg(dino/diffusion/10) 98.6 99.0 86.6 63.7 87.0
Diff-Reg(dino/backboneepnp) 95.8 96.1 88.7 69.0 87.4
Implementation Details. For the single-pass backbone design, we follow 2D3D-MATR [4]. Specifically, for images in data pair, we utilize a ResNet [15] with FPN [68] to generate down-sampled superpixel and associated features. For the point cloud in the data pair, we exploit KPConv[14] to extract the down-sampled superpoint and associated features. A transformer[54] is deployed with inputs of superpixel and associated image features (including ResNet features and DINO features) from the image and superpoint and associated features from the target point cloud to predict the cross-modality features. The denoising transformer in $g_{\theta}$ design has a similar definition. The new non-metric point cloud generated from the image depth map of the Depth anything model[52] are used as inputs of the weighted SVD part in $g_{\theta}$ to compute $R,t$. We utilize the coarse level circle loss[4] and fine level matching loss[4] for the singe-pass backbone[4] and a focal loss for our denoising module $g_{\theta}$. The framework is trained and evaluated with PyTorch on one NVIDIA RTX 3090 GPU. We train our model about 30 epochs with batch size 1.
The overview of our framework for the 2D3D registration task is illustrated in Fig.5. We use Diff-Reg(dino) to denote the single-pass prediction head trained solely. Diff-Reg(dino/backbone) refers to the single-pass prediction head jointly trained with our denoising module $g_{\theta}$. Diff-Reg(dino/diffusion/#) denotes that the ”Patch Matching” module in the top half (in Fig.5) is replaced with our denoising module (with # steps of reverse sampling). The coarse level image features $F^{\hat{X}}$, $F^{\hat{X}}_{dino}$, and point features $F^{\hat{Q}}$ are treated as const inputs of denoising module $g_{\theta}$ in each reverse sampling step. For all variants of Diff-Reg(*), we exclude the three scales ”6 × 8, 12 × 16, 24 × 32” (refer to section 4.1 in [4]) after the coarsest level of ResNet and retain only ”24 × 32”. The visual features from DINOv2 [39] are combined with the coarsest level features that are upsampled to the finest level in ResNet. The training and reverse sampling process are listed in Algorithm.4 and Algorithm.5.
Denoising Module $g_{\theta}$ for 2D3D Registration. In this section, we elucidate the functionality of our denoising module (see Fig.5). By leveraging the recent SOTA depth estimation model DepthAnything [52], we consider this non-metric depth as a constant input for each image. Diff-Reg(dino/backboneepnp) (in Table.V) utilizes the EPnP solver [55] to compute the transformation $R,t$. Its high recall provides evidence that the implicit data augmentation from our diffusion model in matrix space indeed enhances the image backbone and the point cloud backbone. However, the EPnP solver’s vulnerability to outliers during inference hinders its direct application in the reverse sampling process. Integrating the more robust PnP solver [69] into our denoising module could potentially remove the need for the depth estimation model. This improvement is an aspect we aim to explore in future research.
During inference, each reverse sampling step takes $E^{t-1}$ from the previous step to provide initial correspondences. We then use either weighted SVD or the EPnP solver to calculate the transformation $T^{t-1}$ that maps the point cloud $\hat{Q}$ into the image plane space. Subsequently, we apply $T^{t-1}$ to $\hat{Q}$ to get warped $\hat{Q}^{T^{t-1}}$. Next, we use $\hat{Q}^{T^{t-1}}$, $\hat{X}$, $F^{\hat{X}}_{dino}$, $F^{\hat{X}}$, and $F^{\hat{Q}}$ as inputs of the transformer in $g_{\theta}$ to compute the updated image features and point features for calculating the pseudo ground truth $\hat{E}^{0}$.
Figure 5: Overview of our diffusion matching model framework for 2D3D registration task: The top half (the gray part framed by the dashed line) is the single-pass prediction head with the backbones, while the bottom is our denoising module. The transformer in our denoising module utilizes the $F^{\hat{X}}_{dino}$, $F^{\hat{X}}$, $F^{\hat{Q}}$, $\hat{X}$, $\hat{Q}$, and $D^{\hat{X}}$ as inputs, while ”patch matching” module in the single-pass head takes $F^{\hat{X}}_{dino}$, $F^{\hat{X}}$, $F^{\hat{Q}}$, $\hat{X}$, and $\hat{Q}$ as inputs. Please zoom in for details.
(g) Diff-Reg(dino)
(h) Diff-Reg(dino/diffusion/10)
@@
Algorithm 4 Training Diff-Reg in 2D-3D Registration Task
1:Coarse level image and points $\hat{\mathbf{X}}\in\mathbb{R}^{\hat{H}{\times}{\hat{W}}{\times}{3}}$, $\hat{\mathbf{Q}}\in\mathbb{R}^{3}$ and associated features $\mathbf{F}^{\hat{\mathbf{X}}},\mathbf{F}^{\hat{\mathbf{X}}}_{dino},\mathbf{F}^{\hat{\mathbf{Q}}}$.
2:while not converged do
3:     Sample $\mathbf{E}^{0}\sim q(\mathbf{E}^{0})$
4:     $N\times M\leftarrow\mathbf{E}^{0}.\text{shape}$
5:     Sample $t\sim\text{Uniform}(1,...,T)$
6:     $\epsilon\sim\mathcal{N}(0,1)^{N\times M}$
7:     $\mathbf{E}^{t}\leftarrow\sqrt{\bar{\alpha}_{t}}\mathbf{E}^{0}+\sqrt{1-\bar{\alpha}_{t}}f_{\epsilon}(\epsilon),\quad\tilde{\mathbf{E}}^{t}=\mathbf{E}^{t}-\rm{Min}(\mathbf{E}^{t})$
8:     $\hat{\mathbf{E}}_{0}\leftarrow g_{\theta}(\tilde{\mathbf{E}}^{t},\hat{\mathbf{X}},\hat{\mathbf{Q}},\mathbf{F}^{\hat{\mathbf{X}}},\mathbf{F}^{\hat{\mathbf{X}}}_{dino},\mathbf{F}^{\hat{\mathbf{Q}}})$
9:     Optimize $L_{t}=\text{Focal\_loss}(\hat{\mathbf{E}}_{0},\mathbf{E}^{0})$
10:end while
Algorithm 5 Sampling by Diff-Reg in 2D-3D Registration Task
1:Initial matching matrix $\mathbf{E}^{T}$ from backbone or white noise; Coarse level image and points $\hat{\mathbf{X}}\in\mathbb{R}^{\hat{H}{\times}{\hat{W}}{\times}{3}}$, $\hat{\mathbf{Q}}\in\mathbb{R}^{3}$ and associated features $\mathbf{F}^{\hat{\mathbf{X}}},\mathbf{F}^{\hat{\mathbf{X}}}_{dino},\mathbf{F}^{\hat{\mathbf{Q}}}$.
2:Target matching matrix $\tilde{\mathbf{E}}^{0})$.
3:$N\times M\leftarrow\mathbf{E}^{T}.\text{shape}$
4:for $t=T,...,1$ do
5:     ${\mathbf{z}_{t}}\sim\mathcal{N}(0,1)^{N\times M}$ if $t>1$ else ${\mathbf{z}_{t}}\leftarrow\mathbf{0}^{N\times M}$
6:     $\hat{\mathbf{E}}_{0}\leftarrow g_{\theta}(\tilde{\mathbf{E}}^{t},\hat{\mathbf{X}},\hat{\mathbf{Q}},\mathbf{F}^{\hat{\mathbf{X}}},\mathbf{F}^{\hat{\mathbf{X}}}_{dino},\mathbf{F}^{\hat{\mathbf{Q}}})$
7:     $\epsilon_{t}\leftarrow\frac{\hat{\mathbf{E}}_{0}}{\sqrt{1-\bar{\alpha}_{t}}}-\frac{\sqrt{\bar{\alpha}_{t}}}{\sqrt{1-\bar{\alpha}_{t}}}\hat{\mathbf{E}}^{t}$
8:     $\sigma_{t}\leftarrow\sqrt{\frac{(1-\alpha_{t-1})}{1-\alpha_{t}}}\sqrt{1-\frac{\alpha_{t}}{\alpha_{t-1}}}$
9:     $\hat{\mathbf{E}}^{t-1}\leftarrow\sqrt{\alpha_{t-1}}\hat{\mathbf{E}}_{0}+\sqrt{1-\alpha_{t-1}-\sigma^{2}_{t}}\epsilon_{t}+\sigma_{t}\mathbf{z}_{t}$
10:end for
11:$\tilde{\mathbf{E}}^{0}=\mathbf{f}_{\rm{sinkhorn}}(\tilde{\mathbf{E}}^{0})$
Datasets. RGB-D Scenes V2[70] is generated from 14 indoor scenes with 11427 RGB-D frames. Following[4], we split 14 sequences of image-to-point-clouds pairs data, where scenes $0-8/9-10/11-14$ for training/validation/testing. The resulting dataset contains $1748/236/497$ image-to-point-cloud pairs.
[…]
Quantitative results. Following [4], we compare our diffusion matching model with FCGF [65], P2-Net [37], Predator [11], 2D3D-MATR [4], and FreeReg [38]. We introduce a robust single-pass baseline “Diff-Reg(dino)”, in which we integrate the visual feature backbone DINOv2 [39] into the single-pass model 2D3D-MATR [4]. “Diff-Reg(dino/backbone)” represents the single-pass prediction head derived from “Diff-Reg(dino)” after joint training with our denoising module $g_{\theta}$. “Diff-Reg(Diffusion/1)” and “Diff-Reg(Diffusion/10)” refer to our diffusion matching model with one step and ten steps of reverse denoising sampling.
As demonstrated in Table V, the results for “Diff-Reg(dino/backbone)” indicate that the diffused training samples in the matrix space serve as data augmentation to enhance the representation of the ResNet feature backbone and single-pass prediction head. Additionally, the outcome for “Diff-Reg(dino/diffusion/10)” illustrates that our denoising module $g_{\theta}$ effectively tackles the scale ambiguous issue in 2D3D registration. Furthermore, the result for “Diff-Reg(dino/diffusion/1)” reveals that the diffused training samples within the matrix space indeed enhance the single-pass prediction head in “Diff-Reg(dino)”.
We also carried out an additional experiment to show that the performance improvement of “Diff-Reg(dino/backbone)” compared to “Diff-Reg(dino)” is attributed to our diffusion matching model. In this experiment, we did not use the depth map and instead employed one differentiable EPnP [67] solver in the denoising module $g_{\theta}$ to replace the weighted SVD layer. We inputted only superpixels and superpoints with weights computed by associated superpixel and superpoint features. The single-pass prediction head “Diff-Reg(dino/backboneepnp)” (in Table V) achieved a high recall rate of 87.4%, indicating that our diffused samples in matching matrix space indeed enhance the feature backbone.
@@
Qualitative results. The examples of ”Diff-Reg(dino/diffusion/10)” in Fig. 6 reveal that our diffusion matching model excels at capturing salient correspondences crucial for combinatorial consistency, regardless of their distance from the camera. On the other hand, the matches generated by “Diff-Reg(dino)” tend to be more focused at specific distances (from camera). For instance, in the third row, the correspondences produced by “Diff-Reg(dino)” are located very close to the camera, while the correspondences from “Diff-Reg(dino/diffusion/10)” encompass objects such as hats on the black table that are situated at a greater distance. In the first row, “Diff-Reg(dino)” fails to capture the correspondences on the sofa, and in the second row, the correspondences of the white hat on the table are lost. An extreme case in the fourth row demonstrates that “Diff-Reg(dino)” misses the farthest correspondence on the column bookshelf or wall. We provide more visualizations in Fig. 8 to illustrate our method’s effectiveness.
(a) GeoTR[6]
(b) RoITr[2]
(c) Diff-Reg(Diffusion/20)
Figure 7: More qualitative results of non-rigid registration in the 4DMatch/4DLoMatch benchmark. The blue and yellow colors denote the source and target point cloud, respectively. The green and red lines indicate whether the threshold accepts the predicted deformable flow from the source points. The deformable registration is built by GraphSCNet[59]. Zoom in for details.
(g) Diff-Reg(dino)
(h) Diff-Reg(dino/diffusion/10)
Figure 8: More qualitative results on the RGB-D Scenes V2 benchmark[70]. The green/red color indicates whether the matching score is accepted based on a threshold value. Zoom in for details.
IV-C Ablay Study and Discussion
Capability of Escaping from Local Minima. The DDPM framework is specifically designed for removing noise from perturbed samples. Typically, we begin with a baseline method that provides an initial solution, which can then be further refined to achieve better performance. To demonstrate that our denoising network has indeed learned the posterior distribution for the denoising module, we conducted an ablative experiment. In this experiment, we initiated reverse sampling from the solution obtained by the single-pass prediction head or from Gaussian white noise. In Table VI, the results of $E^{T}_{Backbone}$ demonstrate that our denoising network can overcome local minima produced by the single-pass prediction head.
TABLE VI: $E^{T}_{Backbone}$ denote the starting point where $E^{T}$ is generated by the single-pass prediction head. $E^{T}_{Gaussian}$ denote the starting point where $E^{T}$ is sampling from the Gaussian white noise $N(0,1)^{N{\times}M}$. $z_{t}=0$ denotes deterministic sampling, while $z_{t}\neq 0$ denotes the random sampling.
$z_{t}\neq 0$ $z_{t}=0$
3DMatch 3DLoMatch 3DMatch 3DLoMatch
FMR IR RR FMR IR RR FMR IR RR FMR IR RR
$E^{T}_{Backbone}$ 96.12 31.07 94.8 69.79 9.71 73.8 96.30 30.92 94.9 69.71 9.51 73.4
$E^{T}_{Gaussian}$ 96.14 31.07 94.6 69.73 9.72 73.8 96.28 30.93 94.8 69.62 9.56 73.3
$z_{t}\neq 0$ $z_{t}=0$
4DMatch 4DLoMatch 4DMatch 4DLoMatch
NFMR IR NFMR IR NFMR IR NFMR IR
$E^{T}_{Backbone}$ 88.38 86.38 75.94 67.64 88.34 86.36 76.22 67.82
$E^{T}_{Gaussian}$ 88.40 86.40 76.09 67.73 88.72 86.72 76.48 68.16
Reverse Sampling Steps. In our approach, we consider the denoising module as an optimizer that searches for the optimal matching matrix. We argue that increasing the number of search iterations may lead to better solutions. To validate this hypothesis, we experimented to investigate the impact of iterative searching steps on the results. We ran the reverse sampling step from 1 to 20 iterations. As shown in Table. VIII, the registration recall of our diffusion matching model increases as the number of sampling steps grows. The results prove that reverse sampling can reach a better solution by increasing the number of search steps.
TABLE VII: Comparison of time cost in the reverse sampling step. The best results are highlighted in bold.
Iters 3DMatch 3DLoMatch
RR Time(sec.) RR Time(sec.)
GeoTR [6] 92.0 0.296 74.0 0.284
GeoTR. + PEAL [33] 1-step 93.7 0.663 77.8 0.642
KPConv. + Diff-Reg 1-step 94.8 0.048 73.2 0.052
GeoTR. + PEAL [33] 5-step 94.0 2.131 78.5 2.074
KPConv. + Diff-Reg 5-step 94.8 0.182 73.7 0.194
TABLE VIII: The ablay study of the iterative steps of reverse sampling.
Iters RR NFMR/IR
3DMatch 3DLoMatch 4DMatch 4DLoMatch
1 94.8 73.3 85.34/83.93 73.11/65.26
10 94.8 73.4 87.99/86.07 75.46/67.15
20 95.0 73.8 88.39/86.64 76.16/67.82
The lightweight design of our denoising module. Due to the lightweight design, our denoising network showcases a notable speed enhancement compared to other concurrent diffusion-based methods for rigid point cloud registration. This efficiency improvement allows our approach to complete more denoising iterations. We have included a detailed list displaying the time costs (refer to Table VII) associated with each reverse denoising step, along with comparisons to recent studies. Our method achieves competitive results on the 3DMatch benchmark while maintaining fast processing speeds.
(a) GeoTR[6]
(b) RoITr[2]
(c) Diff-Reg(Diffusion/20)
Figure 9: Failure cases of our diffusion matching model on 4DMatch/4DLoMatch benchmark. The blue and yellow colors denote the source and target point cloud, respectively. The green and red lines indicate whether the threshold accepts the predicted deformable flow from the source points. The deformable registration is built by GraphSCNet[59]. Zoom in for details.
V Challenges and Limitations
As illustrated in Fig. 9, deformable registration encounters various challenges, including large motion, low overlap, sparse overlapping regions, and local non-rigid motion (e.g., the dress worn by the dancer in the last line of Fig. 9). Although our diffusion matching model is intended to address both rigid and deformable registration tasks, it may have inherent limitations. The transformer in our denoising module, along with the associated KPConv and ResNet backbone, follows a relatively generic design. To overcome these challenges, we acknowledge the importance of integrating more robust feature embedding techniques [6, 2] into our feature backbone for enhancements. Furthermore, incorporating task-specific priors into our denoising module could prove beneficial in addressing specific issues such as local non-rigid motion. We would leave these for future work.
The success of rigid registration with low overlapping relies on discriminative local feature embedding. The suboptimal performance of our method on the 3DLoMatch benchmark can be linked to the lack of integrated geometric embeddings, such as those mentioned in [34, 5, 2, 6], within feature backbone or transformer in our denoising module. We aim to address this as an area for improvement in future research.
VI Conclusion

### v4

@@ 4 Experiments
Qualitative Results. We provide a visualization to demonstrate our method’s effectiveness in Fig.2. For a fair comparison, we exploit the source point cloud’s “metric index” (i.e., the test point set in the 4DMatch/4DLoMatch dataset) for all methods. Taking the predicted correspondences from RoITr [66], GeoTr [66], and “Our (steps=20)” as anchor correspondences, we calculate the deformation flow for the source test points by applying neighborhood k-nearest neighbors (KNN) interpolation based on the anchors. The deformable registration of the bear’s two front paws in the first row and second/fourth column reveals that dealing with ambiguous matching patches of asymmetric objects can be highly challenging. However, our denoising process can handle this scenario perfectly.
The deformable registration results from the top three rows (from the 4DMatch benchmark) indicate that the baseline methods struggle to compute reliable and consistent correspondences between scans with large deformations. The bottom two rows (from the 4DLoMatch benchmark) also demonstrate that low overlapping combined with deformation results in a disaster. These visualizations demonstrate that our denoising module in the matching matrix space provides a more effective approach for tackling deformable registration tasks.
4.2 2D-3D Registration Task: RGB-D SCENES V2
The 2D-3D registration task is non-trivial because the 2D image data is a perspective projection of the 3D scene, which creates the scale ambiguity problem [25]. Since there is no good robust differentiable PnP [23] solver that can be integrated into our denoising module, we deploy the SOTA depth estimation model DepthAnything [61] to generate “affine-invariant depth” for the image. Subsequently, we utilize the camera’s intrinsic matrix to project the estimated depth map onto a new non-metric point cloud paired with the corresponding real point cloud in the dataset, creating a challenge scale ambiguous registration problem. To alleviate the image data’s scale ambiguity, we use the SOTA self-supposed pre-trained visual feature backbone DINOv2 [36] to enhance the image features.
Datasets. RGB-D Scenes V2 [22] are generated from 14 indoor scenes comprising 11,427 RGB-D frames. Following [25], we split the 14 sequences into image-to-point-cloud pairs data, where scenes 0-8/11-14/9-10 are used for training/validation/testing. The resulting dataset contains 1,748 training pairs, 236 validation pairs, and 497 testing pairs of image-to-point-clouds.
4.2.1 Implementation Details.
For the single-pass backbone design, we follow 2D3D-MATR [25]. Specifically, for images in data pair, we utilize a ResNet [17] with FPN [31] to generate down-sampled superpixel and associated features. For the real point cloud in the data pair, we exploit KPConv [45] to extract the down-sampled superpoints with associated features. A transformer [47] is deployed with inputs of superpixel and associated image features (including ResNet features and DINO features) from the image and superpoint and associated features from real point cloud to predict the cross-modality features. The denoising transformer in $g_{\theta}$ design has a similar definition. The new non-metric point cloud generated from the image depth map output by the DepthAnything model [61] is used as inputs of the weighted SVD function in $g_{\theta}$ to compute $\mathbf{R},\mathbf{t}$.
We utilize the coarse level circle loss [25] and fine level matching loss [25] for the singe-pass backbone [25], while a focal loss for our denoising module $g_{\theta}$. More details about network design are in the appendix. We train our model about 30 epochs with batch size 1.
Metrics. We evaluate our method using the Registration Recall (RR) metric: the ratio of image-to-point-cloud pairs’ RMSE is under 10cm.
@@
Table 3: Evaluation results on RGB-D Scenes V2 [25]. The best results are highlighted in bold, and the second-best results are underlined.
Method Scene-11 Scene-12 Scene-13 Scene-14 Mean
@@
Registration Recall(%$\uparrow$
@@
Diff-Reg(dino/steps=1) 94.4 98.0 85.6 63.7 85.4
Diff-Reg(dino/steps=10) 98.6 99.0 86.6 63.7 87.0
@@
Quantitative Results. We introduce a single-pass baseline “Diff-Reg(dino)”, in which we integrate the visual foundation model DINOv2 [36] into the single-pass model 2D3D-MATR’s [25] ResNet. “Diff-Reg(dino/backbone)” represents the single-pass prediction head derived from “Diff-Reg(dino)” after joint training with our denoising module $g_{\theta}$. “Diff-Reg(steps=1)” and “Diff-Reg(steps=10)” refer to our diffusion matching model with one step and ten steps of reverse denoising sampling.
As demonstrated in Table 3, the results for “Diff-Reg(dino/backbone)” indicate that the diffused training samples in the matrix space serve as data augmentation to enhance the representation of the ResNet feature backbone and single-pass prediction head. Additionally, the outcome for “Diff-Reg(dino/steps=10)” reveals that our denoising module $g_{\theta}$ effectively tackles the scale ambiguous issue in the 2D-3D registration. Furthermore, the result for “Diff-Reg(dino/steps=1)” reveals that the diffused training samples within the matrix space indeed enhance the single-pass prediction head in “Diff-Reg(dino).”
We also carried out an additional experiment to show that the performance improvement of “Diff-Reg(dino/backbone)” compared to “Diff-Reg(dino)” is attributed to our diffusion matching model. In this experiment, we did not use the depth map (from DepthAnything model[61]), opting instead to employ a differentiable weighted EPnP [23] solver in the denoising module $g_{\theta}$ to replace the weighted SVD layer. We take only superpixels and superpoints as inputs of the EPnP solver, with correspondence weights computed by associated superpixel and superpoint features. This setting denoted by “Diff-Reg(dino/backboneepnp)” (in Table 3) achieved a high recall rate of 87.4%, indicating that our diffused samples in matching matrix space indeed enhance the feature backbone.
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
@@
Qualitative Results. The prediction examples of "Diff-Reg(dino/steps=10)" in Fig. 3 reveal that our diffusion matching model excels at capturing salient correspondences crucial for combinatorial consistency, regardless of their distance from the camera. On the other hand, the matches generated by “Diff-Reg(dino)” tend to be more focused at specific distances. For instance, in the third row, the correspondences produced by “Diff-Reg(dino)” are located very close to the camera, while the correspondences from “Diff-Reg(dino/steps=10)” encompass objects such as hats on the black table that are situated at a greater distance. In the first row, “Diff-Reg(dino)” fails to capture the correspondences on the sofa, and in the second row, the correspondences of the white hat on the table are lost. An extreme case in the fourth row demonstrates that “Diff-Reg(dino)” misses the farthest correspondence on the column bookshelf or wall.
5 Conclusion